# Qffusion: Controllable Portrait Video Editing via Quadrant-Grid Attention Learning

### v1

@@ 5.3 Quantitative Comparison on Animation
We report PSNR, SSIM, Warp Error, and LPIPS for quantitative comparison in Table 2. Our Qffusion excels the current state-of-the-art methods by a large margin on PSNR, SSIM, and LPIPS. Although GAN can achieve a slightly superior Warp Error than our Qffusion, it yields the worst ID fidelity in Fig. 4. Besides, Qffusion achieves the best Warp Error among diffusion-based methods, which shows our capacity for long-term temporal modeling by Quadrant-grid Arrangement (QGA) and Quadrant-grid Propagation (QGP). Besides, although the temporal module of AnimateDiff can bring temporal prior, it only supports fixed and limited length generation. Even if the overlap strategy is used for long video generation, the Warp Error of ControlNet $+$ AnimateDiff is still the worst.
@@
5.4 Ablation Studies
Figure 7: We compare with three recent video editing methods: TokenFlow [16], Codef [39] and AnyV2V [27] on portrait video editing. TokenFlow cannot deal with certain local editing (e.g., hair editing), Codef always has artifacts, and AnyV2V struggles for a pleasing appearance. Our method requires both modified start and end frames ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) as editing signals, where $\mathbf{I}^{e}$ is omitted here for simplicity.
Figure 8: Our Qffusion can realize fine-grained local editing, including changing age and style, and adding beauty masks. Note that Qffusion needs the modified start and end frame ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) for editing propagation, where $\mathbf{I}^{e}$ is omitted for simplicity.
@@ 5.4 Ablation Studies
Quadrant-grid Propagation. To validate the effectiveness of our QGP inference, we compare it with the naive inference in Tab. 4. QGP outperforms the naive inference on all metrics, especially on Warp Error, which decreases 72.6% dramatically. The excellent Warp Error of QGP shows its superiority in long-term temporal consistency, which achieves intervals between each sub-image more even in our quadrant-grid design. Besides, we provide the qualitative comparison between the proposed QGP inference and the naive inference in Fig. 6, which also support our findings in Tab. 4.
@@ 5.5 Applications
Portrait video editing. We compare Qffusion with the current state-of-the-art video editing methods Codef [39], TokenFlow [16] and AnyV2V [27] in Fig. 7. The editing scenarios consist of modifying style and hair, and adding sunglasses. As a text-driven editing method, TokenFlow [16] cannot deal with some fine-grained local editing, such as hair editing. Codef [39] sometimes suffers from the inconsistent ID issue (left example). Note that Codef [39] relies on generating a canonical image to record static content in a video, which however usually are with artifacts. When we further edit it with the desired style or appearance, these artifacts will also be spread throughout the entire video. In contrast, our editing results are identity-consistent and abide by the conditional poses clearly. Moreover, although AnyV2V [27] performs image-driven video editing, it faces the degradation of editing appearance. Further, none of the existing methods can generate arbitrary-long videos.
We also provide more examples of our Qffusion on portrait video editing in Fig. 8, where more intermediate frames are given. To further evaluate the video editing ability of Qffusion, we provide a quantitative comparison in Tab. 5. Specifically, we report the average CLIP-Image similarity and Warp Error. Our method yields the best performance on CLIP score and Warp Error, which demonstrates that Qffusion can achieve amazing reference alignment and motion consistency. Note that we cannot calculate CLIP-Image for TokenFlow [16], since it is a text-driven method. Besides, one may wonder whether we can use image editing tools [7, 59] for portrait video editing. We argue that per-frame editing often fails to preserve temporal consistency. Thus, we do not compare with per-frame editing results.
[…]
TokenFlow [16] - 0.874
Codef [39] 0.938 0.301
AnyV2V [27] 0.924 0.822
Ours 0.959 0.206
Table 5: Quantitative comparison with other video editing techniques on CLIP-Image similarity and Warp Error.
Whole-body driving. In Fig. 9, we display our Qffusion can also perform whole-body animation, where UBCFasion [58] dataset is applied. For the conditions, we use DWPose [55] to detect landmarks for the face and body. Specifically, we provide the visual comparison with the current state-of-the-art methods: DreamPose [25], BDMM [57], and Animate Anyone [22]. We apply the re-produced version [23] for Animate Anyone since its official code and dataset are not publicly available. These methods are either carefully designed for whole-body animation tasks (i.e., DreamPose, BDMM), or need an additional heavy appearance network to encode the appearance of input identities (i.e., Animate Anyone). It is difficult for DreamPose to ensure clothing consistency. Besides, Animate Anyone struggles to guarantee facial fidelity. Without additional modules, our method can obtain better results than these task-specific methods, demonstrating its generalizability.
@@
In addition to the qualitative comparison, we also give the quantitative comparison in Tab. 6. It can be seen that our method can achieve better results than those task-specific state-of-the-art techniques.
@@
Figure 11: User study on the selected ratio. Our Qffusion outperforms the current state-of-the-art video editing methods.
@@ 6 Conclusion
To ensure the reproducibility and completeness of this paper, we make the Appendix with 2 sections. Appendix A provides more details for our QGP inference pipeline. Appendix B presents more of our Qffusion’s video animation capacity. Appendix C discusses the limitations of our method.
@@ Appendix A More Details for QGP Inference
We give an example of our QGP inference strategy. Assuming there is an 82-frame video indexed as $\{0-th,1-st,...,80-th,81-st\}$ to be generated, in the first iteration, we apply the $0-th$ and $81-st$ frames as references to generate two intermediate frames: $27-th$ and $54-th$. Then, in the second iteration, we use the newly generated $27-th$ and $54-th$ as references. Concretely, we make $\{0-th,27-th\}$, $\{27-th,54-th\}$ and $\{54-th,81-st\}$ as references separately to synthesize new intermediate frames, forming $\{0-th,9-th,18-th,27-th\}$, $\{27-th,36-th,45-th,54-th\}$, and $\{54-th,63-th,72-nd,81-st\}$. The process would be stopped until all frames are produced. Note that our QGP inference gradually uses generated frames at the current inference iteration (e.g., $27-th$ and $54-th$) as reference frames for the next iteration, making the intervals between each sub-image in the four-grid representation the same in each iteration.
In contrast, the naive inference method fixs the start and end frames ($0-th$ and $81-st$) as references to gradually generate all intermediate frames $\{1-st,2-nd,...79-th,80-th\}$.
Figure 12: More qualitative comparisons with other video animation methods.
@@ Appendix B More Results on Portrait Video Animation
Recall that in Fig. 4 of our main paper, we provide a qualitative comparison with different video animation methods, including: (i) GAN [32], (ii) ControlNet [60], and (iii) ControlNet $+$ Animatediff [18] . Here, we provide more visual comparison in Ap-Fig. 12. The results demonstrate that compared with these animation methods, our Qffusion performs the best animation in terms of ID consistency and condition alignment.
Appendix C Limitations
While our method provides promising results for various applications, there still exist some limitations. For instance, as seen in Fig. 13, our Qffusion sometimes faces unsatisfying cross-ID portrait video animation results. The reason here is that when the driving landmarks come from a different person, the shape information cannot be well-aligned between different persons.

### v2

@@ 5.3 Quantitative Comparison on Animation
We report PSNR, SSIM, Warp Error, and LPIPS for quantitative comparison in Tab. 2. Our Qffusion excels the current state-of-the-art methods by a large margin on PSNR, SSIM, and LPIPS. Although GAN can achieve a slightly superior Warp Error than our Qffusion, it yields the worst ID fidelity in Fig. 4. Besides, Qffusion achieves the best Warp Error among diffusion-based methods, which shows our capacity for long-term temporal modeling by Quadrant-grid Arrangement (QGA) and Quadrant-grid Propagation (QGP). Besides, although the temporal module of AnimateDiff can bring a temporal prior, it only supports fixed and limited length generation. Even if the overlap strategy is used for long video generation, the Warp Error of ControlNet $+$ AnimateDiff is still the worst. Further, the performance of Animate Anyone* and MagicAnimate are also inferior to our method on portrait video animation.
5.4 Ablation Studies
@@
Quadrant-grid Propagation. To validate the effectiveness of our QGP inference, we compare it with the naive inference in Tab. 4. QGP outperforms the naive inference on all metrics, especially on Warp Error, which decreases 72.6% dramatically. The excellent Warp Error of QGP shows its superiority in long-term temporal consistency, which achieves intervals between each sub-image more even in our quadrant-grid design. Besides, we provide the qualitative comparison between the proposed QGP inference and the naive inference in Fig. 6, which also supports our findings in Tab. 4.
Figure 7: We compare Qffusion with four recent video editing methods: TokenFlow [19], Rerender-A-Video [64], Codef [44] and AnyV2V [31] on portrait video editing. Besides, we provide animation results from ControlNext-SVD [46]. Note that our method requires both modified start and end frames ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) as editing signals, where $\mathbf{I}^{e}$ is omitted here for simplicity.
@@
[…]
TokenFlow [19] - 0.874
Rerender-A-Video [64] - 0.759
Codef [44] 0.938 0.301
ControlNext-SVD [46] 0.875 3.606
AnyV2V [31] 0.924 0.822
Ours 0.959 0.206
Table 5: Quantitative comparison with other video editing techniques on CLIP-Image similarity and Warp Error. Besides, we provide the performance of ControlNext-SVD [46].
@@ 5.5 Applications
Portrait video editing. We compare Qffusion with the current state-of-the-art video editing methods Codef [44], Rerender-A-Video [64], TokenFlow [19] and AnyV2V [31] in Fig. 7. The editing scenarios consist of modifying style and hair, and adding sunglasses. As a text-driven editing method, TokenFlow [19] and Rerender-A-Video [64] cannot deal with some fine-grained local editing, such as hair editing. Codef [44] sometimes suffers from the inconsistent ID issue (left example). Note that Codef [44] relies on generating a canonical image to record static content in a video, which, however, usually has artifacts. When we further edit it with the desired style or appearance, these artifacts will also be spread throughout the entire video. In contrast, our editing results are identity-consistent and abide by the conditional poses clearly. Moreover, although AnyV2V [31] performs image-driven video editing, it faces the degradation of editing appearance. Besides, we use ControlNext-SVD [46] for comparison, which integrates ControlNet [71] into an I2V model SVD [5] for video controllable video generation. Here, we use facial keypoints as driving conditions. However, it yields inferior performance to our Qffusion. More importantly, none of the existing methods can generate arbitrary-long videos.
@@
We also provide more examples of our Qffusion on portrait video editing in our Appendix. To further evaluate the video editing ability of Qffusion, we provide a quantitative comparison in Tab. 5. Specifically, we report the average CLIP-Image similarity and Warp Error. Our method yields the best performance on CLIP score and Warp Error, which demonstrates that Qffusion can achieve amazing reference alignment and motion consistency. Note that we cannot calculate CLIP-Image for TokenFlow [19] and Rerender-A-Video [64], since they are text-driven methods.
Whole-body driving. In Fig. 8, we display our Qffusion can also perform whole-body animation, where UBCFasion [69] dataset is applied. For the conditions, we use DWPose [65] to detect landmarks for the face and body. Specifically, we provide the visual comparison with the current state-of-the-art methods: DreamPose [29], BDMM [67], and Animate Anyone [26]. We apply the re-produced version [27] for Animate Anyone since its official code and dataset are not publicly available. These methods are either carefully designed for whole-body animation tasks (i.e., DreamPose, BDMM), or need an additional heavy appearance network to encode the appearance of input identities (i.e., Animate Anyone). It is difficult for DreamPose to ensure clothing consistency. Besides, Animate Anyone struggles to guarantee facial fidelity. Without additional modules, our method can obtain better results than these task-specific methods, demonstrating the generalizability of the proposed method.
Figure 10: User study on the selected ratio. Our Qffusion outperforms the current state-of-the-art video editing methods in all three aspects.
We also give the quantitative comparison in Tab. 6. It can be seen that our method can achieve better results than those task-specific state-of-the-art techniques. The corresponding video results are shown in our supplement.
@@ 5.6 User Study
6 Limitations
While our method provides promising results for various applications, there still exist some limitations. For instance, as seen in Fig. 11, Qffusion sometimes faces unsatisfying cross-ID portrait video animation results. The reason here is that when the driving landmarks come from a different person, the shape information cannot be well-aligned.
@@ 7 Conclusion
To ensure the reproducibility and completeness of this paper, we make this Appendix with 4 sections. Appendix A provides more details for our QGP inference pipeline. Appendix B presents more of our portrait video animation capacity. Appendix C shows that we can perform arbitrary-long video animation and editing. Appendix D shows more results on portrait video editing.
@@ Appendix A More Details for QGP Inference
Ap-Fig. 1: More qualitative comparisons with other video animation methods using explicit landmarks.
Ap-Fig. 2: Qualitative and Quantitative comparisons with LivePortrait [21], which is a non-diffusion animation method with implicit keypoints rather than explicit landmarks. Our Qffusion can achieve competitive or better results with LivePortrait.
Ap-Fig. 3: Qffusion can generate arbitrary-long videos while maintaining appearance and motion consistency, where we show 502-frame results. Top: using the naive and QGP inference to perform long-term video animation. Bottom: using the naive and QGP inference to perform long-term video editing.
Ap-Fig. 4: Quantitative comparison between the naive and QGP inference for long-term video generation. Here, we use PSNR and SSIM for illustration.
We give an example of our QGP inference strategy. Assuming there is an 82-frame video indexed as $\{0-th,1-st,...,80-th,81-st\}$ to be generated. Then, in the first iteration, we apply the $0-th$ and $81-st$ frames as references to generate two intermediate frames: $27-th$ and $54-th$. Next, in the second iteration, we use the newly generated $27-th$ and $54-th$ as references. Concretely, we make $\{0-th,27-th\}$, $\{27-th,54-th\}$ and $\{54-th,81-st\}$ as references separately to synthesize new intermediate frames, forming $\{0-th,9-th,18-th,27-th\}$, $\{27-th,36-th,45-th,54-th\}$, and $\{54-th,63-th,72-nd,81-st\}$. The process would be stopped until all frames are produced. Note that our QGP inference gradually uses generated frames at the current inference iteration (e.g., $27-th$ and $54-th$) as reference frames for the next iteration, making the intervals between each sub-image in the four-grid representation the same in each iteration.
In contrast, the naive inference method fixes the start and end frames ($0-th$ and $81-st$) as references to gradually generate all intermediate frames $\{\{1-st,80-th\},\{2-nd,79-th\},...\}$.
Ap-Fig. 5: Our Qffusion can realize fine-grained local editing, including changing age and style, and adding beauty masks. Note that Qffusion needs the modified start and end frame ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) for editing propagation, where $\mathbf{I}^{e}$ is omitted for simplicity.
Ap-Fig. 6: When the edited start and end frames are inconsistent, our method can present interesting transition results. We give two sets of examples, in which the start and end frames are slightly inconsistent and completely inconsistent, respectively.
Ap-Fig. 7: Warp Error and CLIP-Image similarity across consistency conditions. Here, we calculate CLIP-Image similarity from the edited start frame and end frame, respectively.
@@ Appendix B More Results on Portrait Video Animation
Recall that in Fig. 4 of our main paper, we provide a qualitative comparison with different video animation methods, including: (i) GAN [37], (ii) ControlNet [71], (iii) ControlNet $+$ Animatediff [22], (iv) Animate Anyone* [27], and (v) MagicAnimate [63]. Here, we provide more visual comparison in Ap-Fig. 1. The results demonstrate that compared with these animation methods, our Qffusion performs the best portrait-video animation on ID consistency and condition alignment.
In addition to the above explicit-landmark-based animation methods, recent LivePortrait [21] explores implicit-keypoint-based non-diffusion framework. Besides, LivePortrait and our Qffusion have the following differences. (i) The former uses a two-stage and mixed image-video training strategy, while our Qffusion can be trained in an end-to-end manner. (ii) Our Qffusion can easily be extended to whole-body video, while LivePortrait cannot. As seen in Ap-Fig. 2, we provide a qualitative and quantitative comparison with LivePortrait. Our Qffusion can achieve competitive or better results with LivePortrait.
Appendix C Long Video Animation and Editing
As a recursive inference method, our Quadrant-grid Propagation (QGP) also accumulates errors over time. However, it is worth noting that compared with the naive inference method, our QGP alleviates the error-accumulation issue by making a more even temporal sampling.
For qualitative illustration, we provide a 502-frame animation and editing results using the naive and QGP inference in Ap-Fig. 3, which shows our capability of long-term stability. Besides, in Ap-Fig. 4, we provide a quantitative comparison between the naive and QGP inferences using PSNR and SSIM. It can be seen that the naive inference method accumulates exaggerated errors for long-term videos. It first uses $0-th$ and $501-th$ frames as reference frames to generate two frames $\{1-st,500-th\}$. Then it uses these two generated frames $\{1-st,500-th\}$ as references to generate the $\{2-nd,499-th\}$ frames. That is, the process is conducted iteratively, where the newly generated frame pair $\{l-th,(501-l)-th\}$ is used as reference to synthesize the next frame pair $\{(l+1)-th,(500-l-1)-th\}$, until all intermediate frames are synthesized. Due to error accumulation, the later the generated frame, the worse the quality.
In contrast, our QGP inference has the same interval between sub-images of each four-grid representation. Specifically, in the first iteration, we apply the $0-th$ and $501-st$ frames as references to generate two intermediate frames: $167-th$ and $334-th$. Next, in the second iteration, we make $\{0-th,167-th\}$, $\{167-th,334-th\}$ and $\{334-th,501-st\}$ as references separately to synthesize new intermediate frames. The process would be stopped until all frames are produced. From Ap-Fig. 4, we find that our QGP inference can greatly alleviate error accumulation even in long-term video generation.
Appendix D More Results on Portrait Video Editing
To demonstrate that our Qffusion can generate videos aligning with the given conditions smoothly, we also provide more portrait video editing examples in Ap-Fig. 5, where more driving conditions are presented.
Besides, one may wonder whether our method can deal with potential inconsistencies between the edited first and last frames. We give two sets of examples in Ap-Fig. 6, in which the edited start and end frames are slightly inconsistent and completely inconsistent, respectively. Our Qffusion can present smooth transition results. Besides, in Ap-Fig. 7, we quantitatively measure CLIP-Image similarity and Warp Error of editing video under three cases: ”completely consistent”, ”slightly inconsistent” and ”completely inconsistent”. Here, CLIP-Image score is calculated by the edited start and end video frames, respectively. We found that when the edited first and last frames are inconsistent, the generated video tends to quickly align with the appearance of the end frame, which makes the CLIP-Image score of the edited end frame higher than that of the edited start frame. Besides, Warp Error gradually increases when the edited start and end frames are gradually inconsistent.